A password for the computer is a string of digits and/or letters. The password must obey the following rules:
a. The length of the password must be exactly 8 characters
b. The password must contain at least one digit, and no more than three digits.
What is the number of different passwords you can form?
1
Expert's answer
2020-05-21T17:56:56-0400

we have 10 options for digit - 0-9 and 26 for letter - a-z, so

N=10*267+102266+103*265=80318101760+30891577600+11881376000=123091055360